Sarah has two rectangular sheets of metal that measure 14 inches (in.) by 22 in. She formed Box A by cutting 2 in. squares from each corner of the sheet of metal and folding up the flaps to form a rectangular prism with an open top. Sarah then formed Box B by cutting 1 in. squares from the corners of the other sheet of metal and folding up the flaps to form another rectangular prism with an open top. What is the difference in the volumes, in cubic inches, of Box A and Box B?

Answer by ankor@dixie-net.com(22740)

Sarah has two rectangular sheets of metal that measure 14 inches (in.) by 22 in.
She formed Box A by cutting 2 in. squares from each corner of the sheet of metal and folding up the flaps to form a rectangular prism with an open top.
Sarah then formed Box B by cutting 1 in. squares from the corners of the other sheet of metal and folding up the flaps to form another rectangular prism with an open top.
What is the difference in the volumes, in cubic inches, of Box A and Box B?
:
Box A dimensions; (2" squares subtracts 4" from the length and width)
14 - 4 = 10" width
22 - 4 = 18" length
2 in is the height
10 * 18 * 2 = 360 cu/in
:
Box B dimensions; (1" squares subtracts 2")
14 - 2 = 12"
22 - 2 = 20
1 in is height
12 * 20 * 1 = 240 cu/in
:
360 - 240 = 120 cu/in difference